OpenStudy (anonymous):

Let f be the function from R to R defined by f(x) = x2• Find

OpenStudy (anonymous):

\[F(x)=x ^{2}\]

OpenStudy (anonymous):

\[a. f ^{-1}\left( \left\{ 1 \right\} \right)\]

OpenStudy (anonymous):

\[b. f ^{-1}\left( \left\{ x, 0<x<1 \right\} \right)\]

OpenStudy (anonymous):

\[c. f ^{-1}\left( \left\{ x,x>4 \right\} \right)\]

OpenStudy (anonymous):

u want to evaluate inverse function in general i think ???

OpenStudy (anonymous):

please give me the answer I dont know how to evaluate it

OpenStudy (anonymous):

\[f(x)=x^2\]inverse of this function :\[f^{-1}(x)=\sqrt{x}\]

OpenStudy (anonymous):

oh, tks